USING DIFFUSION MONTE CARLO TO PROBE ROTATIONAL EXCITED STATES

\underline{ANDREW S. PETIT} and ANNE B. McCOY, Department of Chemistry, The Ohio State University, Columbus, OH 43210.

Since its inception in 1975 by Anderson,^{a} Diffusion Monte Carlo (DMC) has been successfully applied to a wide range of electronic and vibrational problems. In the latter case, it has been shown to be a powerful method for studying highly fluxional systems exhibiting large amplitude vibrational motions. We report here our recent work developing a new DMC algorithm capable of treating rotational excited states. We first develop the appropriate coordinates, nodal structures, and re-crossing corrections for this problem. Then, using H_{3}O^{+} and D_{3}O^{+} as model systems,^{b} we show that our method can successfully describe a range of rotational states from |0, 0, 0⟩ to \frac{1}{\sqrt{2}} (|10, 10, 0⟩+|10, −10, 0⟩). In particular, we examine the combined effects of rotational and zero-point vibrational motion on the geometric structure of the molecules. Finally, we find the |10, 0, 0⟩ state to be somewhat problematic but show that the problem is straightforward to identify and has a well-defined solution.

aJ. B. Anderson, J. Chem. Phys., \underline{63}, 1499 (1975).

bX. Huang, S. Carter, and J. Bowman, J. Chem. Phys., \underline{118}, 5431 (2003).